# Transferable Reinforcement Learning via Generalized Occupancy Models

### v2

@@ 7 Discussion
This work introduced generalized occupancy models, a method for transferable reinforcement learning that does not incur compounding error or test-time policy optimization. GOMs can quickly adapt to provide optimal policies for any reward by modeling the distribution of all possible future outcomes along with policies to reach them. We presented an efficient algorithm to learn GOMs and demonstrated the benefits of GOMs over standard successor features and model-based RL techniques. The limitations of our work open future research opportunities. First, GOMs require a choice of features $\phi(s)$ that linearly express the rewards; this assumption may fail, necessitating more expressive feature learning methods. Second, GOMs model the behavior distribution of the dataset; hence, policy optimality can be affected by dataset skewness, which motivates the use of more efficient exploration methods for data collection. Finally, the current version of GOMs infer the reward from offline state-reward pairs; a potential future direction could apply this paradigm to online adaptation, where the reward is inferred from online interactions.
@@
“Transferable Reinforcement Learning via Generalized Occupancy Models”
@@ Appendix A Missing Proofs
We simplify the planning phase of GOM into the following form: for a given reward weight $w_{r}$, in each time step, we have
@@
Appendix B Alternative Derivation of Guided Diffusion Planner
In this section, we derive the guided diffusion planner from the perspective of control as inference [32]. To start, we define the trajectory-level optimality variable $\mathcal{O}$ as a Bernoulli variable taking the value of 1 with probability $\exp(R(\tau))$ and 0 otherwise, where $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r(s_{t})-R_{\text{max}}$. Note we subtract the max discounted return $R_{\text{max}}$ to make the density a valid probability distribution. Planning can be cast as an inference problem where the goal is to sample $\psi^{*}\sim p(\psi|O)$. By Bayes rule, we have
@@ C.2 Training details
For each environment, we train our models on the offline dataset for 100,000 gradient steps using the AdamW optimizer [34] with batch size 2048. The learning rate for the outcome model and the policy are set to $3e^{-4}$ and adjusted according to a cosine learning rate schedule with 500 warmup steps. We train the diffusion noise predictor with 1000 diffusion timesteps. To generate samples from a distribution, we maintain an exponential moving average model with decay rate $0.995$ and employ the DDIM sampler with 50 timesteps [46] on the EMA models. For reward identification, we perform stochastic gradient descent on the regression weights $w$ with learning rate $3e^{-4}$. The same set of training hyperparameters is shared across all evaluation domains.
We use the guided diffusion planner for transfer. The guidance coefficient $\alpha$ is $0.5$ for antmaze, $0.01$ for Franka kitchen, and $0.05$ for Roboverse. We found planning to be sensitive to the guidance coefficient. Hence, for new environments, we suggest using the random shooting planner to get a baseline performance and then tuning the guided diffusion coefficient to improve performance and speed up inference. We run each experiment with 4 random seeds and report the mean and standard deviation in the tables. Each experiment (pretraining + adaptation) takes 3 hours on a single Nvidia L40 GPU.
@@ C.3 Baselines
We use the implementation of Universal Successor Features [5] and Forward-Backward Representation [53, 54] from the code release of [54]. We choose random Fourier features for universal SF as it performs best across the evaluation suite. To ensure fairness of comparison, we set the feature dimension to be 128 for both Universal SF and FB. Both methods are pretrained for 1 million gradient steps and adapted using the same relabelled data as GOMs.
@@
We adapt the original RaMP implementation [8] and convert it into an offline method. RaMP originally consists of an offline training and an online adaptation stage, where online adaptation alternates between data collection and linear regression. We instead transfer by relabelling the offline dataset with the test-time reward function, thus removing the exploration challenge. We use an MPC horizon of 15 for all experiments.
@@ Appendix D Environment Details
D4RL Antmaze [15] is a navigation domain that involves controlling an 8-DoF quadruped robot to reach some designated goal location in a maze. Each task corresponds to reaching a different goal location. We use the standard D4RL offline dataset for pretraining. For downstream task adaptation, we replace the standard sparse reward $\mathbbm{1}(s=g)$ with a dense reward $\exp(-||s-g||^{2}_{2}/20)$ to mitigate the challenge of sparse reward in long-horizon problems.
Franka Kitchen Franka Kitchen [17] is a robotics domain where the goal is to control a Franka arm to interact with appliances in the kitchen. Each task corresponds to interacting with a set of items in no particular order. We use the standard D4RL offline dataset for pretraining. For downstream task adaptation, we use the Markovian sparse rewards, where at each timestep the robot gets a reward equal to the number of completed tasks. We report the number of tasks completed throughout the entire episode in Table 1.
Preference Antmaze Preference Antmaze is a variant of D4RL Antmaze [14] where the goal is to reach the top right cell from the bottom left cell in a custom maze shown in Fig. 4. The two tasks in this environment are the two paths to reaching the goal, simulating different human preferences. To construct the dataset, we collect 1 million transitions using the D4RL waypoint controller. For each preference, we design a reward function that encourages the agent to take one path and not the other. Fig. 6 visualizes the dataset and the reward function for each preference.
Roboverse Robotic Manipulation Roboverse [45] is a tabletop manipulation environment consisting of a WidowX arm aiming to complete multi-step problems. Each task consists of two phases, and the offline dataset contains separate trajectories for each phasebut not full task completion. We use the standard sparse reward, assigning a reward of 1 for each timestep the task is completed.
@@ Appendix E Additional Experiments
Table 4: Full offline multitask RL on AntMaze and Kitchen. GOMs show superior transfer performance (in average episodic return) than successor features, model-based RL, and misspecified goal-conditioned baselines, while being competitive with an oracle using privileged information.
GOM (Ours) USF FB RaMP MOPO COMBO GC-IQL GC-Oracle
@@
To understand the impact of various design decisions on the performance of GOMs, we conducted systematic ablations on the various components, using the antmaze-medium-diverse-v2 as a test bed.
@@ E.1 Full D4RL Results
Table. 4 displays the full D4RL results with oracle goal-conditioned baseline labeled GC-Oracle. The baseline is trained on a goal distribution covering the testing-time goals, granting it privileged information. Despite this, GOMs are competitive with the oracle in most domains.
@@ E.3 Ablation of Feature Dimension and Type
To understand the importance of feature dimension and type, we compare variants of our method that use lower-dimensional random Fourier features, plain random features, and the two top-performing pretrained features from [54]. From Table. 7 we observe that as feature dimension decreases, their expressivity diminishes, resulting in lower performance. We found random features to perform much worse than random Fourier features. Interestingly, pretrained features with dynamics prediction and graph Laplacian objectives also achieve lower returns than random Fourier features. We hypothesize these pretrained features overfit to the training objective and are less expressive than random Fourier features
@@ E.4 Ablation of Dataset Coverage
We investigate the effect of dataset coverage on the performance of our method. We compare GOMs trained on the full D4RL dataset against two variants, one where we randomly subsample half of the transitions, and the other where we adversarially remove the transitions from the half of the state space containing the test-time goal. As shown in Table 7, the performance of GOM drops as dataset coverage degrades.
E.5 Nonparametric Baseline
@@
To confirm the intuition of our method, we implemented a nonparametric baseline that constructs an empirical estimate of the outcome distribution. First, we calculate the empirical discounted summ of features along trajectories in the dataset. Given the set of empirical $(s,a,\psi)$ pairs and the downstream reward weight $w$, we can then select the optimal action at state $s$ through the following steps: (1) query the $k$ nearest neighbors of $s$, (2) evaluate their corresponding values $w^{\top}\psi$, (3) take the action of the top-valued neighbor.
We found this baseline to perform surprisingly well on antmaze-medium-diverse-v2. While it does not achieve the performance of GOMs, it outperforms the RaMP and MOPO baselines. This result confirms the intuition behind the GOM design, which involves selecting the optimal outcome under dataset coverage and taking an action to realize it. We attribute the performance gap between this baseline and GOMs to their trajectory stitching ability (acquired via dynamic programming), infinite horizon modeling, and neural network generalization.
@@ Appendix F Algorithm Pseudocode
Algorithm 2 GOM Offline Adaptation
@@
Algorithm 3 GOM Online Adaptation
@@
Algorithm 4 GOM Inference (Random Shooting)
@@
Algorithm 5 GOM Inference (Guided diffusion)

### v3

@@ 7 Discussion
This work introduced Distributional Successor Features for Zero-Shot Policy Optimization (DiSPOs), a method for transferable reinforcement learning that does not incur compounding error or test-time policy optimization. By modeling the distribution of all possible future outcomes along with policies to reach them, DiSPOs can quickly provide optimal policies for any reward in a zero-shot manner. We presented an efficient algorithm to learn DiSPOs and demonstrated the benefits of DiSPOs over standard successor features and model-based RL techniques. The limitations of our work open future research opportunities. First, DiSPOs require a choice of features $\phi(s)$ that linearly express the rewards; this assumption may fail, necessitating more expressive feature learning methods. Second, DiSPOs model the behavior distribution of the dataset; hence, policy optimality can be affected by dataset skewness, which motivates the use of more efficient exploration methods for data collection. Finally, the current version of DiSPOs infer the reward from offline state-reward pairs; a potential future direction could apply this paradigm to online adaptation, where the reward is inferred from online interactions.
Acknowledgment
CZ is supported by the UW-Amazon fellowship. TH is supported by the NSF GRFP under Grant No. DGE 2140004. SSD acknowledges the support of NSF IIS 2110170, NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF CCF 2019844, and NSF IIS 2229881.
@@ Acknowledgment
“Distributional Successor Features Enable Zero-Shot Policy Optimization”
@@ Appendix A Missing Proofs
We simplify the policy optimization phase of DiSPOs into the following form: for a given reward weight $w_{r}$, in each time step, we have
@@
Appendix B Alternative Derivation of Guided Diffusion Sampling
In this section, we derive the guided diffusion sampling from the perspective of control as inference [33]. To start, we define the trajectory-level optimality variable $\mathcal{O}$ as a Bernoulli variable taking the value of 1 with probability $\exp(R(\tau))$ and 0 otherwise, where $R(\tau)=\sum_{t=0}^{T}\gamma^{t}r(s_{t})-R_{\text{max}}$. Note we subtract the max discounted return $R_{\text{max}}$ to make the density a valid probability distribution. Planning can be cast as an inference problem where the goal is to sample $\psi^{*}\sim p(\psi|O)$. By Bayes rule, we have
@@ C.2 Training details
We train our models on the offline dataset for 100,000 gradient steps using the AdamW optimizer [35] with batch size 2048. The learning rate for the outcome model and the policy are set to $3e^{-4}$ and adjusted according to a cosine learning rate schedule with 500 warmup steps. We train the diffusion noise prediction network with 1000 diffusion timesteps and sample using the DDIM [47] sampler with 50 timesteps. We sample from an exponential moving average model with decay rate $0.995$. The same set of training hyperparameters is shared across all environments.
To transfer to a downstream task, we randomly sample 10000 transitions from the dataset, relabel them with the test time rewards, and perform linear least squares regression. We use the guided diffusion planner for Antmaze ($\alpha=0.5$), Franka Kitchen ($\alpha=0.01$), and Roboverse ($\alpha=0.05$) experiments. For Hopper, we use the random shooting planner with 1000 particles. We found planning with guided diffusion to be sensitive to the guidance coefficient. Hence, for new environments, we suggest using the random shooting planner to get a baseline performance and then tuning the guided diffusion coefficient to acclerate inference. We run each experiment with 6 random seeds and report the mean and standard deviation in the tables. Each experiment (pretraining + adaptation) takes 3 hours on a single Nvidia L40 GPU.
@@ C.3 Baselines
We use the implementation of Universal Successor Features Approximators [5] and Forward-Backward Representation [54, 55] from the code release of [55]. We choose random Fourier features for universal SF as it performs best across the evaluation suite. To ensure fairness of comparison, we set the feature dimension to be 128 for both Universal SF and FB. Both methods are pretrained for 1 million gradient steps and adapted using the same reward estimation method as DiSPOs.
@@
We adapt the original RaMP implementation [9] and convert it into an offline method. RaMP originally consists of an offline training and an online adaptation stage, where online adaptation alternates between data collection and linear regression. We instead adapt by subsampling 10000 transitions from the the offline dataset and relabeling them with the test-time reward function, thus removing the exploration challenge. We use an MPC horizon of 15 for all experiments.
@@ Appendix D Environment Details
[16] is a navigation domain that involves controlling an 8-DoF quadruped robot to reach some designated goal location in a maze. Each task corresponds to reaching a different goal location. We use the standard D4RL offline dataset for pretraining. For downstream task adaptation, we replace the standard sparse reward $\mathbbm{1}(s=g)$ with a dense reward $\exp(-||s-g||^{2}_{2}/20)$ to mitigate the challenge of sparse reward in long-horizon problems.
Franka Kitchen [18] is a manipulation domain where the goal is to control a Franka arm to interact with appliances in the kitchen. Each task corresponds to interacting with a set of items in no particular order. We use the standard D4RL offline dataset for pretraining. For downstream task adaptation, we use the Markovian sparse rewards, where at each timestep the robot gets a reward equal to the number of completed tasks. We report the number of tasks completed throughout the entire episode in Table 1.
Hopper Hopper [7, 9] is a locomotion domain where the task is to control a hopper to perform various tasks, including hopping forward, hopping backward, standing, and jumping. We design a shaped reward for each task. We pretain on the offline dataset from [9], collected by taking the replay buffers of expert SAC [19] agents trained to solve each task.
Preference Antmaze is a variant of D4RL Antmaze [15] where the goal is to reach the top right cell from the bottom left cell in a custom maze shown in Fig. 4. The two tasks in this environment are the two paths to reaching the goal, simulating different human preferences. To construct the dataset, we collect 1 million transitions using the D4RL waypoint controller. For each preference, we design a reward function that encourages the agent to take one path and not the other. Fig. 6 visualizes the dataset and the reward function for each preference.
Roboverse [46] is a tabletop manipulation environment consisting of a WidowX arm aiming to complete multi-step problems. Each task consists of two phases, and the offline dataset contains separate trajectories for each phasebut not full task completion. We use the standard sparse reward, assigning a reward of 1 for each timestep the task is completed.
@@ Appendix E Additional Experiments
Table 4: Full offline multitask RL on AntMaze and Kitchen. DiSPOs show superior transfer performance (in average episodic return) than successor features, model-based RL, and misspecified goal-conditioned baselines, while being competitive with an oracle using privileged information.
DiSPO (Ours) USFA FB RaMP MOPO COMBO GC-IQL GC-Oracle
@@
hopper-forward 566 $\pm$ 63 487 $\pm$ 110 452 $\pm$ 59 470 $\pm$ 16 493 $\pm$ 114 982 $\pm$ 157 - -
hopper-backward 367 $\pm$ 15 261 $\pm$ 68 269 $\pm$ 77 220 $\pm$ 15 596 $\pm$ 211 194 $\pm$ 74 - -
hopper-stand 800 $\pm$ 0 685 $\pm$ 130 670 $\pm$ 120 255 $\pm$ 15 800 $\pm$ 0 600 $\pm$ 111 - -
hopper-jump 832 $\pm$ 22 746 $\pm$ 112 726 $\pm$ 35 652 $\pm$ 28 753 $\pm$ 51 670 $\pm$ 109 - -
@@
To understand the impact of various design decisions on the performance of DiSPOs, we conducted systematic ablations on the various components, using the Antmaze medium-diverse as a test bed.
@@ E.1 Full D4RL Results
Table. 4 displays the full D4RL results with oracle goal-conditioned baseline labeled GC-Oracle. The baseline is trained on a goal distribution covering the testing-time goals, granting it privileged information. Despite this, DiSPOs are competitive with the oracle in most domains.
@@ E.2 Ablation of Planning Method
Table 8: Ablation of planning horizon.
umaze-diverse medium-diverse
$\gamma=0.99$ 587 $\pm$ 12 650 $\pm$ 50
$\gamma=0.95$ 597 $\pm$ 16 682 $\pm$ 20
$\gamma=0.9$ 601 $\pm$ 19 553 $\pm$ 36
@@ E.3 Ablation of Feature Dimension and Type
To understand the importance of feature dimension and type, we compare variants of our method that use lower-dimensional random Fourier features, vanilla random features, and the two top-performing pretrained features from [55]. From Table. 8 we observe that as feature dimension decreases, their expressivity diminishes, resulting in lower performance. We found random features to perform much worse than random Fourier features. Interestingly, pretrained features with dynamics prediction and graph Laplacian objectives also achieve lower returns than random Fourier features. We hypothesize these pretrained features overfit to the training objective and are less expressive than random Fourier features
@@ E.4 Ablation of Dataset Coverage
We investigate the effect of dataset coverage on the performance of our method. We compare DiSPOs trained on the full D4RL dataset against two variants, one where we randomly subsample half of the transitions, and the other where we adversarially remove the transitions from the half of the state space containing the test-time goal. As shown in Table 8, the performance of DiSPO drops as dataset coverage degrades.
E.5 Ablation of Planning Horizon
In Fig. 8, we ablate the effective planning horizon of DiSPOs by controlling the discount factor $\gamma$. We found that for shorter horizon tasks (umaze-diverse), the optimality of planned trajectories improves as the planning horizon decreases. However, for long-horizon tasks (medium-diverse), reducing the planning horizon too much incurs cost on global optimality.
E.6 Nonparametric Baseline
@@
To confirm the intuition of our method, we implemented a nonparametric baseline that constructs an empirical estimate of the outcome distribution. First, we calculate the empirical discounted sum of features along trajectories in the dataset. Given the set of empirical $(s,a,\psi)$ pairs and the downstream reward weight $w$, we can then select the optimal action at state $s$ through the following steps: (1) query the $k$ nearest neighbors of $s$, (2) evaluate their corresponding values $w^{\top}\psi$, (3) take the action of the top-valued neighbor.
We found this baseline to perform surprisingly well on antmaze-medium-diverse-v2. While it does not achieve the performance of DiSPOs, it outperforms RaMP and MOPO. This result confirms the intuition behind DiSPOs, which involves selecting the optimal outcome under dataset coverage and taking an action to realize it. We attribute the performance gap between this baseline and DiSPOs to their trajectory stitching ability (acquired via dynamic programming), infinite horizon modeling, and neural network generalization.
E.7 Visualization of Stitching
We visualize the stitched trajectories for roboverse environments in Fig. 7. The dataset only contains trajectories from each phase separately, but DiSPOs can generate full trajectories by stitching the subtrajectories.
Figure 7: Visualization of stitched trajectories for roboverse PickPlace and BlockedDrawer.
@@ Appendix F Algorithm Pseudocode
Algorithm 2 DiSPO Offline Adaptation
@@
Algorithm 3 DiSPO Online Adaptation
@@
Algorithm 4 DiSPO Inference (Random Shooting)
@@
Algorithm 5 DiSPO Inference (Guided diffusion)